IF YOU DILUTE 50ML OF A 25% SOLUTION OF TABLE SALT TO 100ML, WHAT WILL BE THE RESULTANT PERCENTAGE STRENGTH?
Answer by stanbon(75887)

IF YOU DILUTE 50ML OF A 25% SOLUTION OF TABLE SALT TO 100ML, WHAT WILL BE THE RESULTANT PERCENTAGE STRENGTH?
----
Equation:
salt = salt
0.25*50 = x*100
(50/4) = 100x
---
x = (1/2)(1/4)
x = 1/8
x = 12.5%
================
Cheers,
Stan H.